OpenStudy (anonymous):

simplify the expression (-5p^4)

Parth (parthkohli):

Is there anything to simplify?

OpenStudy (anonymous):

i got 20p^4

OpenStudy (anonymous):

sorry typo

OpenStudy (anonymous):

(-5p)^4

OpenStudy (unklerhaukus):

\[(ab)^n=a^nb^n\]

OpenStudy (anonymous):

so I was right

OpenStudy (unklerhaukus):

nope

OpenStudy (unklerhaukus):

\[(-5p)^4=(-5)^4p^4\]

OpenStudy (anonymous):

oh ok